OpenStudy (2yoututu11):

What is the exact area of a circle with a radius of 12 feet? A. 24π square feet B. 36π square feet C. 121π square feet D. 144π square feet

OpenStudy (anonymous):

Area = PI * R^2 R = 12, so R^2 = 144 A = 144 * PI.

OpenStudy (anonymous):

\(\ \huge A= \pi r^2 \). So, \(\ \huge A= \pi \times 12^2 \). That's \(\ \huge 144 \pi feet^2 \).